In 15 years, Maya will be twice as old as David is now. In 15 years David will be as old as Maya will be 10 years from now. How old are they now?
Answer by josmiceli(9817)

Read these carefully
d = David's age now
m = Maya's age now
(1) m + 15 = 2 * d
(2) d + 15 = m + 10
from (1)
d = (m + 15) / 2
substituting this in (2)
(m + 15)/2 + 15 = m + 10
multiply through by 2
m + 15 + 30 = 2 * m + 20
subtract m from both sides
subtract 20 from both sides
m = 25
d = (25 + 15) / 2
d = 20
check
20 = (25 + 15) / 2
OK
25 + 15 = 2 * 20
OK